OpenStudy (anonymous):

M=-5 passing through -4,-2

Nnesha (nnesha):

y = mx+b slope intercept form where m is slope b is y-intercept substitute x and y for (-4,-2) point

OpenStudy (tkhunny):

Several "forms" should be in your head. Slope-Intercept is good, but that is not the information you have been given. Point-Slope is good. (y-k) = m(x-h) Point is (h,k) Slope is m. Substitute away!

Nnesha (nnesha):

with given information you can write both slope -intercept `and` point slope equation it depends on the statement. o^_^o